# Prescriptive Agents based on Rag for Automated Maintenance (PARAM)

### v1

@@ II-C Knowledge Layer: Contextual Retrieval and Synthesis
The knowledge layer implements a sophisticated retrieval architecture that combines several complementary approaches. External knowledge integration incorporates real-time web search capabilities for accessing the most current maintenance practices, coupled with API-based connections to specialized industrial databases and maintenance repositories [35]. Internal knowledge management employs vector embedding techniques to create searchable representations of equipment manuals, historical maintenance cases, and codified expert knowledge [36]. The retrieval process incorporates semantic search capabilities that go beyond simple keyword matching to identify conceptually relevant maintenance procedures, even when expressed using different terminology [32]. This is particularly valuable in industrial settings where documentation may use varying nomenclature for similar components or procedures.
@@ II-D Prescriptive Layer: Intelligent Decision Support
The recommendation generation follows a systematic, multi-stage process designed to ensure comprehensive and practical outputs. The initial stage focuses on anomaly contextualization, where the system evaluates fault severity in conjunction with equipment criticality and operational priorities [26]. This assessment draws upon both the immediate detection outputs and historical maintenance patterns to establish appropriate response urgency. Subsequently, the system generates and evaluates a solution space comprising potential maintenance interventions, considering the full spectrum from immediate corrective actions to longer-term preventive strategies [27]. Each potential solution undergoes risk assessment and resource requirement analysis to determine feasibility and impact [37].
@@
Case Fault Type Severity Confidence Findings and Actions
@@ III-A Context Engineering: The Foundation of Intelligent Industrial Maintenance
Context engineering has emerged as a critical discipline in the development of intelligent industrial maintenance systems, fundamentally shifting the paradigm from traditional prompt engineering to comprehensive information orchestration [31]. In the context of the PARAM framework, context engineering serves as the architectural backbone that enables seamless integration of multi-modal data streams, historical maintenance records, and real-time sensor information into actionable maintenance intelligence [26].
The evolution from prompt engineering to context engineering represents more than a terminological distinction—it embodies a fundamental shift in how we approach AI system design [28]. While prompt engineering focused on crafting optimal instructions for individual interactions, context engineering addresses the holistic information ecosystem surrounding AI models, ensuring they have access to relevant, timely, and accurate information throughout complex maintenance workflows [27]. This distinction is particularly crucial in industrial maintenance scenarios where decisions must be grounded in multiple data sources including equipment history, manufacturer specifications, regulatory requirements, and operational constraints [40].
@@ III-B Dynamic Context Assembly in Industrial Maintenance
The dynamic nature of industrial maintenance environments necessitates sophisticated context assembly mechanisms that can adapt to changing operational conditions and equipment states [29]. The PARAM framework’s implementation of context engineering addresses several key challenges identified in the literature [23]. First, the serialization of numerical sensor data into natural language formats enables LLMs to process traditionally numerical domains while maintaining temporal relationships and frequency characteristics critical for bearing fault analysis [30]. This approach bridges the gap between structured numerical data and the text-based processing capabilities of modern language models [32].
The multi-layered context architecture employed in PARAM demonstrates how context engineering can be structured to handle diverse information types [37]. The system maintains:
@@
  • Conversation history for continuity
  • External information retrieval through RAG mechanisms
@@ III-C Knowledge Retrieval and Semantic Integration
The integration of Retrieval-Augmented Generation (RAG) techniques within the PARAM framework exemplifies advanced context engineering principles applied to industrial maintenance [25]. The system’s ability to dynamically retrieve relevant maintenance procedures addresses a fundamental limitation of static knowledge bases [34]. The semantic search capabilities enable the system to identify conceptually relevant maintenance procedures even when expressed using different terminology [35]. The quality and relevance of retrieved context directly impact maintenance recommendation accuracy. Research has demonstrated that careful context selection can improve tool selection accuracy by up to 300% compared to approaches that simply include all available information. This finding is particularly relevant to the PARAM framework’s multi-agentic approach, where different agents must coordinate access to shared knowledge bases while maintaining contextual coherence across the maintenance workflow.
III-D Context Engineering Challenges in Multi-Agent Systems
The implementation of multi-agent systems in industrial maintenance introduces unique context engineering challenges that require sophisticated coordination mechanisms. The literature identifies several critical failure modes that can occur when agents operate without proper context sharing: context confusion, where agents select inappropriate tools or procedures; context poisoning, where erroneous information propagates through the system; and context distraction, where excessive information overwhelms decision-making capabilities.
The PARAM framework’s approach to these challenges through structured context protocols demonstrates best practices for multi-agent coordination. The system implements context isolation techniques that allow specialized agents to operate with focused information sets while maintaining overall system coherence. This approach aligns with emerging protocols such as the Model Context Protocol (MCP) and Agent-to-Agent (A2A) communication standards, which provide standardized mechanisms for context sharing across distributed agent systems.
III-E Small Language Models: The Pragmatic Choice for Industrial Agents
@@
The integration of Small Language Models (SLMs) into the PARAM framework represents a strategic departure from the prevailing emphasis on Large Language Models [39]. Research demonstrates that SLMs, when properly fine-tuned for specific domains, can achieve performance levels comparable to or exceeding larger models on targeted tasks [33].
III-E1 Specialized Task Performance and Domain Adaptation
@@
Research demonstrates that SLMs, when properly fine-tuned for specific domains, can achieve performance levels comparable to or exceeding larger models on targeted tasks. In the context of industrial maintenance, this translates to more reliable fault classification, more accurate severity assessments, and more contextually appropriate maintenance recommendations. The ability to fine-tune SLMs on domain-specific datasets enables the creation of maintenance agents that understand specialized terminology, recognize industry-specific patterns, and generate recommendations aligned with operational constraints.
@@
The modular nature of SLM deployment enables the creation of specialized agent architectures where different models handle different aspects of the maintenance workflow. This approach allows for the development of expert models for specific tasks such as vibration analysis, maintenance scheduling, or parts procurement, each optimized for its particular domain while contributing to the overall maintenance intelligence system. The reduced parameter count of SLMs facilitates rapid iteration and adaptation, enabling maintenance systems to evolve quickly in response to changing operational requirements or emerging failure modes.
III-E2 Resource Efficiency and Edge Deployment
@@
The deployment of SLMs in industrial maintenance environments offers significant advantages in terms of resource utilization and operational independence. The reduced computational requirements of SLMs enable deployment on edge devices and industrial computing platforms without the need for cloud connectivity, addressing critical concerns about latency, bandwidth, and data security in industrial environments. This capability is particularly important for maintenance systems that must operate in environments with limited connectivity or where data sovereignty requirements prevent cloud-based processing.
The energy efficiency of SLMs contributes to sustainable AI deployment in industrial settings. Research indicates that SLM-based systems can achieve operational cost reductions of up to 90% compared to LLM-based approaches while maintaining acceptable performance levels. This cost efficiency enables broader deployment of AI-powered maintenance systems across industrial facilities, democratizing access to advanced maintenance intelligence regardless of organizational size or computational resources.
III-E3 Heterogeneous Agent Architectures
The strategic deployment of SLMs within heterogeneous agent architectures represents an optimal approach to industrial maintenance system design. Rather than relying exclusively on either small or large models, the most effective systems combine the specialized efficiency of SLMs with the general reasoning capabilities of LLMs when necessary. This hybrid approach enables the creation of maintenance systems that default to efficient SLM processing for routine tasks while escalating to more powerful models only when complex reasoning or novel problem-solving is required.
The implementation of dynamic model selection mechanisms allows maintenance systems to automatically route tasks to the most appropriate model based on complexity, urgency, and resource availability. This approach optimizes both performance and resource utilization while maintaining the flexibility to handle exceptional situations that may require more sophisticated reasoning capabilities.
III-F Context Engineering and SLM Integration: Synergistic Benefits
@@
The faster inference times of SLMs enable more interactive context engineering approaches, where the system can engage in multi-turn conversations with maintenance personnel to refine context and improve recommendation quality. This capability is particularly valuable in industrial maintenance scenarios where the initial problem description may be incomplete or where maintenance personnel need to explore multiple solution pathways.
III-G Future Directions and Emerging Challenges
@@
III-H Implications for Industrial Practice

### v2

@@ II-C Knowledge Layer: Contextual Retrieval and Synthesis
The knowledge layer implements a sophisticated retrieval architecture that combines several complementary approaches. External knowledge integration incorporates real-time web search capabilities for access the most current maintenance practices, coupled with API-based connections to specialized industrial databases and maintenance repositories [35]. Internal knowledge management employs vector embedding techniques to create searchable representations of equipment manuals, historical maintenance cases, and codified expert knowledge [36]. The retrieval process incorporates semantic search capabilities that go beyond simple keyword matching to identify conceptually relevant maintenance procedures, even when expressed using different terminology [32]. This is particularly valuable in industrial settings where documentation may use varying nomenclature for similar components or procedures.
@@ II-D Prescriptive Layer: Intelligent Decision Support
The recommendation generation follows a systematic, multi-stage process designed to ensure comprehensive and practical outputs. The initial stage focuses on contextualizing the anomaly, where the system evaluates fault severity in conjunction with equipment criticality and operational priorities [26]. This assessment draws upon both the immediate detection outputs and historical maintenance patterns to establish appropriate response urgency. Subsequently, the system generates and evaluates a solution space comprising potential maintenance interventions, considering the full spectrum from immediate corrective actions to longer-term preventive strategies [27]. Each potential solution undergoes risk assessment and resource requirement analysis to determine feasibility and impact [37].
@@
Case Fault Type Severity Confidence LLM Recommendations
@@ III-A Context Engineering: The Foundation of Intelligent Industrial Maintenance
Industrial machinery comes in various sizes and types, and each manufacturing unit or company has its own set of repair and maintenance manuals. Generic guides often fall short, leading to wasted time, resources, and unresolved issues. However, machine-specific manuals contain detailed solutions for common problems related to that particular model. Identifying the machine type and model, technicians can quickly access the most relevant information for efficient troubleshooting.
Context engineering has emerged as a critical discipline in the development of intelligent systems, in the context of the PARAM framework, context engineering serves as the architectural backbone that enables the seamless integration of multi-modal data streams, historical maintenance records, and real-time sensor information into actionable maintenance intelligence [26].
An LLM-as-a-judge framework was used to evaluate the recommendations generated by different models. The scoring criteria considered the depth of answers, the relevance of recommendations to the breakdown scenario, and the inclusion of specific actionable items rather than generic solutions. In the results, Small Language Models (SLMs) consistently scored lower, which can be attributed to their shorter context window lengths and limitations in retrieved context, reducing their ability to capture the full scope of the problem. Furthermore, their limited capacity to fully interpret complex scenarios often leads them to provide safer, more generalized solutions. However, this trend can be addressed through fine-tuning with domain-specific data, enabling SLMs to deliver more targeted and precise recommendations.
@@ III-B Dynamic Context Assembly in Industrial Maintenance
The dynamic nature of industrial maintenance environments requires sophisticated context assembly mechanisms that can adapt to changing operating conditions and equipment states [29]. The implementation of context engineering of the PARAM framework addresses several key challenges identified in the literature [23]. First, serialization of numerical sensor data into natural language formats enables LLMs to process traditionally numerical domains while maintaining temporal relationships and frequency characteristics critical for bearing fault analysis [30]. This approach bridges the gap between structured numerical data and the text-based processing capabilities of modern language models [32].
The multi-layered context architecture employed in PARAM demonstrates how context engineering can be structured to handle various types of information [37]. The system maintains the following.
@@
  • Conversation history for informed prescriptions on breakdowns
  • External information retrieval through RAG and Web search agent mechanisms
@@ III-C Knowledge Retrieval and Semantic Integration
While our RAG implementation successfully retrieves relevant document chunks, a critical consideration emerges: maintenance prescriptions often require synthesizing information scattered across multiple sections of a manual. In such cases, standard chunking can be a bottleneck, as the critical context is fragmented. Our analysis of context window utilization (typically 3-4% for LLaMA models) confirms the window size itself is not the primary constraint. The quality and relevance of the retrieved context directly impact maintenance recommendation accuracy. Empirical studies on context window utilization in RAG show that optimizing chunk size and the fraction of the context window actually used improves answer quality, and that simply adding more context can degrade performance.[42]
To address this challenge of retrieving a cohesive set of information, a promising strategy is to pre-process maintenance manuals by clustering text not by physical proximity, but by semantic similarity of resolved issues. This creates clusters where all information pertaining to a specific breakdown type—even if originally from different chapters—is grouped, enabling the retrieval of a single, comprehensive context that spans the entire document and provides all necessary details for an accurate, holistic prescription. This is an additional information pre-processing step, which we plan to perform in our future experiments.
@@
The integration of Retrieval-Augmented Generation (RAG) techniques within the PARAM framework exemplifies advanced context engineering principles applied to industrial maintenance [25]. The system’s ability to dynamically retrieve relevant maintenance procedures addresses a fundamental limitation of static knowledge bases [34]. The semantic search capabilities enable the system to identify conceptually relevant maintenance procedures even when expressed using different terminology [35].
III-D Small Language Models: The Pragmatic Choice for Industrial Agents
@@
The integration of Small Language Models (SLMs) into the PARAM framework represents a strategic departure from the prevailing emphasis on Large Language Models (LLMs). Research demonstrates that SLMs, when properly fine-tuned for domain-specific tasks, can achieve performance comparable to or exceeding that of larger models. Our results show that SLMs (LLaMA 3.1 Nano-4B and Nano-8B) are significantly faster than LLMs (e.g., LLaMA 3.1 Ultra 253B). However, as non-fine-tuned models, SLMs exhibit a noticeable drop in accuracy, along with lower completeness and reasoning scores—as evaluated by the LLM Judge benchmark.
III-D1 Specialized Task Performance and Domain Adaptation
@@
Research demonstrates that SLMs, when properly fine-tuned for specific domains, can achieve performance levels comparable to or exceeding larger models on targeted tasks. In the context of industrial maintenance, this results in more reliable fault classification, more accurate severity assessments, and more contextually appropriate maintenance recommendations. The ability to fine-tune SLMs in domain-specific datasets enables the creation of maintenance agents that understand specialized terminology, recognize industry-specific patterns, and generate recommendations aligned with operational constraints.
@@
The modular nature of SLM deployment enables the creation of specialized agent architectures in which different models handle different aspects of the maintenance workflow. This approach allows for the development of expert models for specific tasks such as vibration analysis, maintenance scheduling, or parts procurement, each optimized for its particular domain while contributing to the overall maintenance intelligence system. The reduced parameter count of SLMs facilitates rapid iteration and adaptation, enabling maintenance systems to evolve quickly in response to changing operational requirements or emerging failure modes.
III-D2 Resource Efficiency and Edge Deployment
@@
The deployment of SLMs in industrial maintenance environments offers significant advantages in terms of resource utilization and operational independence. The reduced computational requirements of SLMs enable deployment on edge devices and industrial computing platforms without the need for cloud connectivity, addressing critical concerns about latency, bandwidth, and data security in industrial environments. This capability is particularly important for maintenance systems that must operate in environments with limited connectivity or where data sovereignty requirements prevent cloud-based processing. Evaluations of open-source LLMs in enterprise RAG further indicate that increasing provided context (via higher top‑k) often yields limited answer quality gains, reinforcing the value of focused retrieval and efficient small models for practical deployments [43].
III-E Context Engineering and SLM Integration: Synergistic Benefits
@@
The benchmarking results highlight distinct specializations across models, showing how accuracy, reasoning, cost, and latency trade-offs shape their use cases. Gemini-1.5-Pro (accuracy 4.5, reasoning 4.0, latency 6.5s, overall 4.5) leverages its exceptionally large context window to handle complex, long inputs with strong general-purpose performance, though at higher cost. In contrast, Gemini-1.5-Flash (accuracy 4.5, reasoning 5.0, latency 2.4s, cost \$0.00006, overall 5.0) is optimized for ultra-low latency and efficiency, making it ideal for real-time applications. Building on this, Gemini-2.0-Flash-Exp (accuracy 4.5, reasoning 5.0, latency 2.5s, overall 5.0) sustains the same speed and reasoning excellence. On the open-source side, LLaMA-3.1-Nano-4B (accuracy 3.0, latency 8.49s, overall 3.0) and LLaMA-3.1-Nano-8B (accuracy 3.5, reasoning 3.5, latency 4.66s, overall 3.5) specialize in lightweight, self-hosted deployments where cost efficiency and portability matter more than top-tier reasoning. At the other extreme, LLaMA-3.1-Ultra-253B (accuracy 5.0, reasoning 5.0, latency 83s, overall 4.5) achieves state-of-the-art reasoning and completeness, excelling in high-precision tasks where accuracy outweighs speed and cost. Together, the tableII illustrates a spectrum of trade-offs—from Gemini’s balance of speed and intelligence to LLaMA’s range of language models. Model choice depends on context, constraints and task requirements.
Moreover, the faster inference times of SLMs—such as LLaMA-3.1-Nano-4B with latency 8.49s and overall score 3.0, or LLaMA-3.1-Nano-8B with latency 4.66s and overall score 3.5—enable more interactive context engineering approaches. In practice, higher-end models like Gemini-2.0-Flash-Exp (accuracy 4.5, completeness 5.0, latency 2.5s, overall 5.0) or LLaMA-3.1-Ultra-253B (accuracy 5.0, completeness 5.0, reasoning 5.0, overall 4.5) further highlight how multi-turn conversations with maintenance personnel can refine context and improve recommendation quality. This capability is particularly valuable in industrial maintenance scenarios where the initial problem description may be incomplete or where maintenance personnel need to explore multiple solution pathways.
III-F Future Directions and Emerging Challenges
As observed in our results, Small Language Models (SLMs) score lower than Large Language Models (LLMs). For example, LLaMA-3.1-Nano-4B achieved an accuracy of 3.0, reasoning 3.0, and overall score 3.0 with latency of 8.49 s, while the larger LLaMA-3.1-Ultra-253B reached perfect 5.0 scores in accuracy, completeness, and reasoning, with an overall of 4.5 despite its high latency of 83 s. This demonstrates the natural trade-off between lightweight deployability and advanced reasoning. However, this performance gap can be significantly reduced by fine-tuning the SLMs with machinery-specific datasets—for instance, historical breakdown records and their resolutions, bringing their task-specific performance much closer to LLMs while retaining the benefits of lower inference cost and faster deployment.
@@ III-F Future Directions and Emerging Challenges
III-G Implications for Industrial Practice
The proprietary models (Gemini models) delivered superior results in our tests for instance, Gemini-1.5-Flash achieved an overall score of 5.0 with accuracy 4.5, reasoning 5.0, and a latency of just 2.4 s at a cost of \$0.00006 per query. However, for industrial use cases requiring multiple daily inferences, reliance on proprietary models may not be ideal due to inconsistent responses and high costs. Alternatively, open-source LLMs like LLaMA-3.1 Ultra (253B) showed better accuracy in our experiments, but face challenges with high latency, expensive hosting/inference costs, and inability to deploy on edge devices without cloud interfaces. Our observations also found these open-source LLMs require considerable response time, potentially causing inefficiencies in breakdown maintenance.
In contrast, Small Language Models (SLMs) offer significantly faster response times and easier edge deployment with consistent outputs. However, in our tests, the SLMs were not fine-tuned, resulting in unsatisfactory scores according to the LLM judge evaluation.